\begin{thebibliography}{99}

\bibitem{AhnJFG} M.~W.~Ahn, {\it On the error-sum function of Pierce expansions}, J. Fractal Geom. {\bf 10} (2023), 389--425.
\bibitem{AhnIJNT} M.~W.~Ahn, {\it Convergence exponent of Pierce expansion digit sequences}, Int. J. Number Theory {\bf 21} (2025), 1967--1993.
\bibitem{ERS58} P.~Erd\H os, A.~R\'enyi, and P.~Sz\H usz, {\it On Engel's and Sylvester's series}, Ann. Univ. Sci. Budapest. E\"otv\"os Sect. Math. {\bf 1} (1958), 7--32.
\bibitem{FS24} L.~Fang and L.~Shang, {\it On the exact rate of convergence of digits in Engel expansions}, J. Math. Anal. Appl. {\bf 531} (2024), Paper No. 127726, 16pp.
\bibitem{FW18} L.~Fang and M.~Wu, {\it Hausdorff dimension of certain sets arising in Engel expansions}, Nonlinearity {\bf 31} (2018), 2105--2125.
\bibitem{Gal76} J.~Galambos, {\it Representations of Real Numbers by Infinite Series}, Lecture Notes in Math. 502, Springer, Berlin, 1976.
\bibitem{LW01} Y.-Y.~Liu and J.~Wu, {\it Hausdorff dimensions in Engel expansions}, Acta Arith. {\bf 99} (2001), 79--83.
\bibitem{LW03} Y.-Y.~Liu and J.~Wu, {\it Some exceptional sets in Engel expansions}, Nonlinearity {\bf 16} (2003), 559--566.
\bibitem{LL18} M.~L\"u and J.~Liu, {\it Hausdorff dimensions of some exceptional sets in Engel expansions}, J. Number Theory {\bf 185} (2018), 490--498.
\bibitem{SW20} L.~Shang and M.~Wu, {\it Slow growth rate of the digits in Engel expansions}, Fractals {\bf 28} (2020), 2050047.
\bibitem{SW21} L.~Shang and M.~Wu, {\it On the growth speed of digits in Engel expansions}, J. Number Theory {\bf 219} (2021), 368--385.
\bibitem{She10} L.-M.~Shen, {\it A further discussion of the Hausdorff dimension in Engel expansions}, Acta Arith. {\bf 143} (2010), 271--276.
\bibitem{Wu00} J.~Wu, {\it A problem of Galambos on Engel expansions}, Acta Arith. {\bf 92} (2000), 383--386.

\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage{amsmath,bm,amssymb,amsthm,textcomp}
\usepackage{enumitem}
\usepackage{mathtools}
\usepackage[numbers,sort&compress]{natbib}
\usepackage{hyperref}

\DeclareMathOperator{\id}{id}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}

\numberwithin{equation}{section}

\allowdisplaybreaks

\begin{document}

\title[Topology for Engel expansions: evaluation and coding]{A topology for Engel expansions: evaluation and digit coding maps}

\author{Min Woong Ahn}
\address{Department of Mathematics Education, Silla University, 140, Baegyang-daero 700beon-gil, Sasang-gu, Busan, 46958, Republic of Korea}
\email{minwoong@silla.ac.kr}

\date{\today}

\subjclass[2020]{Primary 37B10; Secondary 11A67, 54E52}
\keywords{Engel expansion, symbolic dynamics, digit coding map, continuity, Baire category}

\maketitle

\begin{abstract}
We develop a topological framework for Engel expansions that treats both directions of the correspondence between points of $(0,1]$ and nondecreasing digit sequences. We endow the sequence space with the product topology to study the evaluation map, and we fix a nonterminating digit algorithm to study the digit coding map. We also record the correspondence between cylinder sets and fundamental intervals, and give an application to Baire category results for functions of the digits.
\end{abstract}

\section{Introduction}

Every $x\in(0,1]$ admits an Engel expansion
\begin{equation}\label{eq:intro-Engel}
x=\frac{1}{d_1(x)}+\frac{1}{d_1(x)d_2(x)}+\frac{1}{d_1(x)d_2(x)d_3(x)}+\dotsb,
\end{equation}
with integer digits $2\le d_1(x)\le d_2(x)\le\dotsb$. Set, for $x\in(0,1]$,
\begin{equation}\label{eq:intro-e1-TE}
d_1(x)\coloneqq\left\lfloor \frac{1}{x}\right\rfloor+1,
\qquad
T(x)\coloneqq d_1(x)x-1\in(0,1],
\end{equation}
and define, for $n\ge1$,
\begin{equation}\label{eq:intro-en}
d_{n+1}(x)=d_1\left(T^{n}(x)\right),
\end{equation}
where $T^n$ denotes the $n$th iterate of $T$. If $1/x\notin\mathbb N$, then $d_1(x)=\lceil 1/x\rceil$ and $0<T(x)<1$; if $x=1/m$ ($m\in\mathbb N$), then $d_1(x)=m+1$ and $T(1/m)=1/m$, so $(d_n(x))_{n\ge1}$ is nondecreasing and eventually constant equal to $m+1$. It is a classical result that the expansion \eqref{eq:intro-Engel} with digits given by \eqref{eq:intro-en} converges to $x$ (see, e.g., \cite{Gal76}).

Two standard, coexisting conventions for rationals are common. One stops the iteration of $x\mapsto x\lceil1/x\rceil-1$ at $0$ and yields a finite expansion. If we adopt the conventions $c\cdot\infty\coloneqq\infty$ ($c>0$) and $1/\infty \coloneqq 0$, then appending an $\infty$-tail to this finite expansion does not affect the resulting sum in \eqref{eq:intro-Engel}. The other---used in \eqref{eq:intro-e1-TE}--\eqref{eq:intro-en}---always replaces $\lceil1/x\rceil$ by $\lfloor1/x\rfloor+1$; at $x=1/m$ it fixes $1/m$ and thus yields the constant tail $m{+}1,m{+}1,\dotsc$ automatically. Both conventions agree on irrationals and differ only at rationals. We adopt the nonterminating convention (without $\infty$-tails) throughout.

Let $\mathbb N_\infty\coloneqq\mathbb N\cup\{\infty\}$, and let $\Sigma_E$ denote the space of Engel digit sequences (nondecreasing sequences in $\mathbb N_\infty$); its precise definition, along with the metric topology, is given in \S\ref{subsec:symbolic-metric}. 
Let $\varphi\colon\Sigma_E\to[0,1]$ and $f\colon(0,1]\to\Sigma_E$ denote the evaluation and digit coding maps, respectively; precise definitions and basic properties are given in \S\ref{subsec:evaluation-coding}.

Fractal aspects of Engel expansions have been widely studied; see, e.g., \cite{Wu00,LW01,LW03,She10,FW18,LL18,SW20,SW21,FS24}. Metric and number–theoretic topics trace back at least to Erd\H os--R\'enyi--Sz\H usz~\cite{ERS58}. Topological uniqueness on irrationals and the behavior at rationals are classical (e.g., \cite{Gal76}); Baire category results for certain level sets were analyzed in \cite{SW21}. A canonical topology on the space of Engel digit sequences, together with a systematic analysis of $\varphi$ and $f$ in that topology, appears to be missing.

We provide such a framework, showing that the evaluation map $\varphi$ is globally Lipschitz continuous (Theorem \ref{thm:phi-cont}) and that the digit coding map $f$ is continuous precisely at the irrationals (Theorem \ref{thm:f-cont-irr}) and one-sided continuous at the rationals (Theorem \ref{thm:f-one-sided}). This clean structural result immediately streamlines Baire category arguments for functions of the digits; in particular, it yields a concise proof that $\{x\in(0,1]:\lambda(x)=\infty\}$, where $\lambda\colon(0,1]\to[0,\infty]$ denotes the {\em convergence exponent} of Engel digit sequences, is comeager in $(0,1]$, thereby recovering \cite[Theorem~3.5]{SW21}. For comparison, an analogous approach for the Pierce expansion (also known as the alternating Engel expansion) appears in \cite{AhnJFG,AhnIJNT}.

This paper is organized as follows. Section~\ref{sec:preliminaries} introduces the symbolic space, the metric, and the associated maps, and recalls the classical fundamental intervals. Section~\ref{sec:main} establishes continuity properties of $\varphi$ and $f$. Section~\ref{sec:applications} applies the framework to show that broad divergence sets contain dense $G_\delta$ subsets (and are therefore comeager), and concludes that $\{x\in(0,1]:\lambda(x)=\infty\}$ is comeager.

\section{Preliminaries}\label{sec:preliminaries}

\subsection{Symbolic space and metric}\label{subsec:symbolic-metric}
Equip $\mathbb N$ with the discrete topology and set $X\coloneqq\mathbb N_\infty=\mathbb N\cup\{\infty\}$. Define
\[
\rho(x,y)\coloneqq
\begin{cases}
0,&x=y,\\
x^{-1}+y^{-1},&x\ne y,
\end{cases}
\qquad (\infty^{-1}\coloneqq0).
\]
By \cite[Lemma~3.1]{AhnJFG}, $\rho$ induces the one-point compactification topology on $X$. On $X^{\mathbb N}$ use the product metric
\begin{equation}\label{eq:metric}
d(\alpha,\beta)\coloneqq\sum_{n=1}^\infty 2^{-n}\rho(\alpha_n,\beta_n),
\qquad\alpha=(\alpha_n)_n,\ \beta=(\beta_n)_n,
\end{equation}
which metrizes the product topology and is topologically equivalent to the factorial-weight metric in \cite[\S3.2]{AhnJFG}.

Define the \emph{space of Engel sequences}
\[
\Sigma_E\coloneqq\left\{\eta=(\eta_n)_{n\ge1}\in X^{\mathbb N}:2\le \eta_1\le \eta_2\le\dotsb\right\},
\]
with the subspace topology from $X^{\mathbb N}$. Put
\[
\Sigma_E^{\mathrm{irr}}\coloneqq\left\{\eta\in \Sigma_E: \eta_n \in \mathbb{N}, \, \forall n \text{ and } \lim_{n\to\infty} \eta_n=\infty\right\},\qquad
\Sigma_E^{\mathrm{rat}}\coloneqq\Sigma_E\setminus\Sigma_E^{\mathrm{irr}}.
\]
Note that $\Sigma_E^{\mathrm{rat}}$ consists of sequences that are eventually constant in $\mathbb{N}$ or attain $\infty$ at some index.

\begin{lemma}\label{lem:SigmaE-closed}
The space of Engel sequences $\Sigma_E$ is closed in $X^{\mathbb N}$.
\end{lemma}

\begin{proof}
If $\beta\notin\Sigma_E$, then either $\beta_1=1$ or there exists $k\ge1$ with $\beta_{k+1}<\beta_k$. In the former case, the cylinder $\{\alpha\in X^{\mathbb{N}}:\alpha_1=1\}$ is an open neighborhood of $\beta$ disjoint from $\Sigma_E$. In the latter case, the cylinder $\{\alpha\in X^{\mathbb{N}}:\alpha_k=\beta_k,\ \alpha_{k+1}=\beta_{k+1}\}$ is open and disjoint from $\Sigma_E$. Hence the complement of $\Sigma_E$ is open.
\end{proof}

\begin{lemma}\label{lem:compact}
The product space $X^{\mathbb N}$ is compact; therefore $\Sigma_E$ is compact.
\end{lemma}

\begin{proof}
Since $X$ is compact (one-point compactification), Tychonoff's theorem implies that $X^{\mathbb N}$ is compact in the product topology, which is metrized by the metric $d$ in \eqref{eq:metric}. Then Lemma~\ref{lem:SigmaE-closed} yields compactness of $\Sigma_E$.
\end{proof}

\subsection{Cylinders and fundamental intervals}\label{subsec:fund-intervals}
A finite sequence $\sigma=(\sigma_1,\dotsc,\sigma_n)$ of integers with $2\le\sigma_1\le\cdots\le\sigma_n$ is called an \emph{admissible word of length $n$} (see \cite[Definition~2.1 \& Proposition~2.2]{SW21}). We write $|\sigma|=n$ for the length of $\sigma$.

For an admissible word $\sigma$ of length $n$, define the \emph{level-$n$ cylinder}
\[
C_n(\sigma)\coloneqq\left\{\eta\in\Sigma_E: \eta_1=\sigma_1,\ \dotsc,\ \eta_n=\sigma_n\right\}.
\]
The corresponding \emph{level-$n$ fundamental interval} is the preimage under $f$:
\[
J_n(\sigma)\coloneqq f^{-1}\left(C_n(\sigma)\right)\subseteq(0,1].
\]

\begin{proposition}[{\cite[p.~84]{Gal76}}]\label{prop:fundamental}
For an admissible word $\sigma$ of length $n$ as above,
\[
J_n(\sigma)=\left(\ell(\sigma),\,r(\sigma)\right],
\]
where
\[
\ell(\sigma)=\sum_{k=1}^n\frac{1}{\sigma_1\dotsm\sigma_k},\quad
r(\sigma)=\sum_{k=1}^{n-1}\frac{1}{\sigma_1\dotsm\sigma_k}
+\frac{1}{\sigma_1\dotsm\sigma_{n-1}(\sigma_n-1)}.
\]
In particular, $C_n(\sigma)$ is clopen in $\Sigma_E$, and $\varphi\left(C_n(\sigma)\right)=[\ell(\sigma),r(\sigma)]$.
\end{proposition}

\begin{remark}
The image $\varphi\left(C_n(\sigma)\right)$ attains its left endpoint $\ell(\sigma)$ by appending $\sigma$ with an $\infty$-tail, $(\sigma_1,\dotsc,\sigma_n,\infty,\infty,\dotsc)$, and its right endpoint $r(\sigma)$ by the largest nonterminating sequence in the cylinder, $(\sigma_1,\dotsc,\sigma_{n-1},\sigma_n,\sigma_n,\dotsc)$ (which is the nonterminating code for $r(\sigma)$).
\end{remark}

\subsection{Evaluation and coding}\label{subsec:evaluation-coding}

The evaluation map $\varphi\colon\Sigma_E\to[0,1]$ is defined by
\[
\varphi(\eta)\coloneqq\sum_{n=1}^\infty\frac{1}{\eta_1\eta_2\dotsm\eta_n}.
\]
Since $(\eta_n)$ is nondecreasing with $\eta_n\ge2$, the tail is dominated by a geometric series, so $\varphi$ is well defined on $\Sigma_E$.
(We adopt the conventions $\infty^{-1}\coloneqq0$ and, for products, that if $\eta_k=\infty$ then $(\eta_1\dotsm\eta_n)^{-1}=0$ for all $n\ge k$; hence the tail vanishes once a coordinate equals $\infty$.)

The digit coding map $f\colon(0,1]\to\Sigma_E$ is given by $f(x)\coloneqq(d_n(x))_{n\ge1}$, where $d_1,\,T,\,d_{n+1}$ are as in \eqref{eq:intro-e1-TE}–\eqref{eq:intro-en}. Under our nonterminating convention, rationals are encoded by eventually constant integer sequences; for instance, $f(3/4)=(2,3,3,3,\dotsc)$. Note that $\varphi(2,3,3,\dotsc)=3/4=\varphi(2,2,\infty,\infty,\dotsc)$, so $\varphi$ is not injective on $\Sigma_E$ at rationals.

\begin{proposition}[\cite{Gal76}]\label{prop:basic-inversion}
We have $\varphi\circ f=\id_{(0,1]}$, and $f\circ\varphi=\id$ on $\Sigma_E^{\mathrm{irr}}$.
\end{proposition}

\begin{proof}
By construction, the Engel expansion of $x\in(0,1]$ is unique, so $\varphi(f(x))=x$.
If $\eta\in\Sigma_E^{\mathrm{irr}}$, then $\eta_n\in\mathbb N$ with $\eta_n\to\infty$, and $\varphi(\eta)$ has Engel digits $\eta_n$; hence $f(\varphi(\eta))=\eta$.
\end{proof}

\section{Main results: continuity and symbolic structure}\label{sec:main}

\subsection{Continuity of the evaluation map}

\begin{theorem}\label{thm:phi-cont}
The evaluation map $\varphi\colon\Sigma_E\to[0,1]$ is Lipschitz (for $d$ in \eqref{eq:metric} and the Euclidean metric).
\end{theorem}

\begin{proof}
Let $\eta,\eta'\in\Sigma_E$ and let $n$ be the first index at which they differ. Put $A\coloneqq1/(\eta_1\cdots\eta_{n-1})=1/(\eta_1'\cdots\eta_{n-1}')$. Then
\[
\varphi(\eta)-\varphi(\eta')
= A\left(\sum_{j=0}^{\infty}\frac{1}{\eta_n\dotsm \eta_{n+j}}
-\sum_{j=0}^{\infty}\frac{1}{\eta_n'\dotsm \eta_{n+j}'}\right).
\]
Since the digits are nondecreasing, $\eta_{n+\ell}\ge \eta_n$ and $\eta_{n+\ell}'\ge\eta_n'$ for all $\ell\ge0$. Hence
\[
\sum_{j=0}^\infty \frac{1}{\eta_n\dotsm\eta_{n+j}} \leq \sum_{j=0}^\infty \eta_n^{-(j+1)}=\frac{1}{\eta_n-1}\le\frac{2}{\eta_n},
\]
and similarly with $\eta$ replaced by $\eta'$. Therefore
\[
\left|\varphi(\eta)-\varphi(\eta')\right|
\le A\left(\frac{2}{\eta_n}+\frac{2}{\eta_n'}\right)
\le 2^{2-n}\left(\frac{1}{\eta_n}+\frac{1}{\eta_n'}\right),
\]
since $\eta_k,\eta_k'\ge2$ imply $A\le 2^{-(n-1)}$. On the other hand,
\[
d(\eta,\eta')\ge 2^{-n}\rho(\eta_n,\eta_n')=2^{-n}\left(\frac{1}{\eta_n}+\frac{1}{\eta_n'}\right).
\]
Combining gives $\left|\varphi(\eta)-\varphi(\eta')\right|\le 4\,d(\eta,\eta')$.
\end{proof}

\subsection{Continuity properties of the coding map}\label{subsec:coding-cont}

Throughout this subsection we work in the compact metric space $(\Sigma_E,d)$ (Lemma~\ref{lem:compact}) and use that $\varphi$ is Lipschitz (Theorem~\ref{thm:phi-cont}). We also recall Proposition~\ref{prop:basic-inversion}: $\varphi\circ f=\id_{(0,1]}$ and $f\circ\varphi=\id$ on $\Sigma_E^{\mathrm{irr}}$.

\begin{lemma}\label{lem:two-codes}
Let $x\in(0,1]\cap\mathbb Q$. Then there is a unique admissible word $\sigma=(\sigma_1,\dots,\sigma_n)$ such that
\[
x=r(\sigma)=\sum_{k=1}^{n-1}\frac{1}{\sigma_1\cdots\sigma_k}
+\frac{1}{\sigma_1\cdots\sigma_{n-1}(\sigma_n-1)}.
\]
Under the nonterminating convention,
\[
\alpha\coloneqq f(x)=(\sigma_1,\dots,\sigma_{n-1},\sigma_n,\sigma_n,\sigma_n,\dotsc).
\]
The terminating Engel code of the same $x$ is obtained by decreasing the last eligible digit by $1$ and appending an $\infty$-tail: there exists an index
\[
j=\max\{1\le k\le n: \sigma_k\ge 3\}
\]
such that
\[
\alpha'\coloneqq(\sigma_1,\dots,\sigma_{j-1},\sigma_j-1,\infty,\infty,\dotsc)\in\Sigma_E
\quad\text{and}\quad
\varphi(\alpha')=\varphi(\alpha)=x.
\]
In particular, $\varphi^{-1}(\{x\})=\{\alpha,\alpha'\}$ and $f(x)=\alpha$.
\end{lemma}

\begin{proof}
The description of $J_n(\sigma)=(\ell(\sigma),r(\sigma)]$ in Proposition~\ref{prop:fundamental} shows that $x$ is the (right) endpoint of a unique fundamental interval, giving the stated $\sigma$ and $\alpha$.
For the terminating code, one checks directly that replacing the rightmost digit $\ge 3$ by $1$ less and appending an $\infty$-tail preserves admissibility and the value under $\varphi$ (the usual carry rule); see also \cite{Gal76}. Uniqueness of the two codes follows from monotonicity and the endpoint structure of fundamental intervals.
\end{proof}

\begin{theorem}\label{thm:f-cont-irr}
The map $f:(0,1]\to\Sigma_E$ is continuous at every irrational $x$.
\end{theorem}

\begin{proof}
Suppose $x\in(0,1]\setminus\mathbb Q$ and $f$ were not continuous at $x$.
Then there are $\varepsilon>0$ and $x_m\to x$ with $d\left(f(x_m),f(x)\right)\ge\varepsilon$ for all $m$.
Since $(\Sigma_E,d)$ is compact (Lemma \ref{lem:compact}), the sequence $(f(x_m))_{m\ge1}$ has a convergent subsequence, say $f(x_{m_k})\to\beta\in\Sigma_E$.
Since $\varphi$ is continuous,
\[
x=\lim_{k\to\infty} x_{m_k}=\lim_{k\to\infty} \varphi\left(f(x_{m_k})\right)=\varphi(\beta).
\]
As $x$ is irrational, $\beta\in\Sigma_E^{\mathrm{irr}}$ and hence
$\beta=f(\varphi(\beta))=f(x)$ (Proposition~\ref{prop:basic-inversion}).
Thus $f(x_{m_k})\to f(x)$, contradicting $d(f(x_{m_k}),f(x))\ge\varepsilon$.
\end{proof}

\begin{theorem}\label{thm:f-one-sided}
Let $x\in(0,1]\cap\mathbb Q$. Then $f$ is left–continuous at $x$ and not right–continuous at $x$.
\end{theorem}

\begin{proof}
Let $\sigma=(\sigma_1,\dots,\sigma_n)$ be as in Lemma~\ref{lem:two-codes}, so
$x=r(\sigma)$ and $J_n(\sigma)=(\ell(\sigma),r(\sigma)]$ with right endpoint $x$.
Write $\alpha=f(x)=(\sigma_1,\dots,\sigma_{n-1},\sigma_n,\sigma_n,\dotsc)$ and
$\alpha'=(\sigma_1,\dots,\sigma_{j-1},\sigma_j-1,\infty,\infty,\dotsc)$ for the terminating code (Lemma~\ref{lem:two-codes}).
Recall from Proposition~\ref{prop:fundamental} that the cylinder
\[
C_n(\sigma)\coloneqq\{\eta\in\Sigma_E:\ \eta_1=\sigma_1,\dots,\eta_n=\sigma_n\}
\]
is \emph{clopen} in $\Sigma_E$.

{\sc Left-continuity.}
Take any sequence $y_m\uparrow x$ with $y_m<x$.
For $m$ large, $y_m\in J_n(\sigma)$, hence $f(y_m)\in C_n(\sigma)$.
Because $C_n(\sigma)$ is closed, any subsequential limit $\gamma$ of $f(y_m)$ lies in $C_n(\sigma)$.
On the other hand, $\varphi$ is continuous, so $\varphi(\gamma)=\lim_m \varphi(f(y_m))=\lim_m y_m=x$.
Thus $\gamma\in\varphi^{-1}(\{x\})=\{\alpha,\alpha'\}$ (Lemma~\ref{lem:two-codes}).
Since $\alpha\in C_n(\sigma)$ but $\alpha'\notin C_n(\sigma)$ (as its first $n$ digits do not match $\sigma$), we must have $\gamma=\alpha=f(x)$.
Therefore $f(y_m)\to f(x)$ as $m\uparrow\infty$.

{\sc Failure of right-continuity.}
Take any sequence $z_m\downarrow x$ with $z_m>x$.
For $m$ large, $z_m\notin J_n(\sigma)$, hence $f(z_m)\notin C_n(\sigma)$.
Since $\Sigma_E\setminus C_n(\sigma)$ is also closed (as $C_n(\sigma)$ is clopen), any subsequential limit $\delta$ of $f(z_m)$ must lie in $\Sigma_E\setminus C_n(\sigma)$.
Again by continuity of $\varphi$,
$\varphi(\delta)=\lim_m \varphi(f(z_m))=\lim_m z_m=x$,
so $\delta\in\{\alpha,\alpha'\}$.
But $\alpha\in C_n(\sigma)$ while $\alpha'\in \Sigma_E\setminus C_n(\sigma)$, hence $\delta=\alpha'\neq f(x)$.
Therefore $\lim_{t\downarrow x} f(t)=\alpha'\ne f(x)$, and $f$ is not right–continuous at $x$.
\end{proof}

\begin{corollary}\label{cor:homeo}
The restriction $\varphi:\Sigma_E^{\mathrm{irr}}\to(0,1]\setminus\mathbb Q$ is a homeomorphism with inverse $f$.
\end{corollary}

\begin{proof}
The bijectivity of $\varphi \colon \Sigma^{\mathrm{irr}}_E \to (0,1] \setminus \mathbb{Q}$ with inverse $f$ is established by Proposition~\ref{prop:basic-inversion}. The continuity of $\varphi$ is a consequence of Theorem~\ref{thm:phi-cont} (Lipschitz continuity). The continuity of the inverse map $f$ on $(0,1]\setminus\mathbb{Q}$ is established by Theorem~\ref{thm:f-cont-irr}. Since $\varphi$ is a continuous bijection with a continuous inverse, it is a homeomorphism.
\end{proof}

\section{Applications: comeager divergence sets and further remarks}\label{sec:applications}

In \cite{AhnIJNT} we studied Pierce digits; here we use the present Engel topology to obtain a general Baire category statement for Engel digits.

\subsection{Comeager divergence for functions of the digits}\label{subsec:comeager-divergence}
For a nonnegative function $\phi\colon\{2,3,\dotsc\}\to[0,\infty)$, write
\[
\mathcal D_\phi\coloneqq\left\{x\in(0,1]: \sum_{n=1}^\infty \phi(d_n(x))=\infty\right\}.
\]

\begin{theorem}\label{thm:Gdelta}
Let $\phi\ge0$ and assume that the set $\{t\ge2: \phi(t)>0\}$ is unbounded. Then there exists a dense $G_\delta$ set $\mathcal G_\phi\subseteq(0,1]$ with $\mathcal G_\phi\subseteq \mathcal D_\phi$. In particular, $\mathcal D_\phi$ is comeager (residual) in $(0,1]$.
\end{theorem}

\begin{proof}
For $m\in\mathbb N$, define
\[
\Upsilon_m\coloneqq
\bigcup_{\substack{\sigma=(\sigma_1,\dotsc,\sigma_n) \text{ admissible}\\ \sum_{k=1}^n \phi(\sigma_k)>m}}
C_n(\sigma) \subseteq \Sigma_E.
\]
Each $\Upsilon_m$ is open (union of clopen cylinders). We claim that $\Upsilon_m\cap\Sigma_E^{\mathrm{irr}}$ is dense in $\Sigma_E^{\mathrm{irr}}$.
Indeed, given a nonempty cylinder $C_N(\tau)$ with $\tau_N\in\mathbb N$ (equivalently, $C_N(\tau)\cap\Sigma_E^{\mathrm{irr}}\neq\varnothing$), choose
$t\ge\max\{2,\tau_N\}$ with $\phi(t)>0$ (possible since $\{t: \phi(t)>0\}$ is unbounded), and append a long enough $t$–tail to $\tau$ so that $\sum_{k=1}^{|\sigma|}\phi(\sigma_k)>m$. Then
$C_{|\sigma|}(\sigma)\subseteq C_N(\tau)\cap\Upsilon_m$, proving density in $\Sigma_E^{\mathrm{irr}}$.

Hence
\[
\Gamma_\phi\coloneqq\bigcap_{m=1}^\infty \Upsilon_m
\]
satisfies that $\Gamma_\phi\cap\Sigma_E^{\mathrm{irr}}$ is a dense $G_\delta$ in $\Sigma_E^{\mathrm{irr}}$; moreover, if $\eta\in\Gamma_\phi\cap\Sigma_E^{\mathrm{irr}}$, then the partial sums $\sum_{k=1}^n\phi(\eta_k)$ are unbounded in $n$, so $\sum_{k\ge1}\phi(\eta_k)=\infty$.

Set
\[
\mathcal G_\phi\coloneqq \varphi\left(\Gamma_\phi\cap\Sigma_E^{\mathrm{irr}}\right)\ \subset\ (0,1]\setminus\mathbb Q.
\]
By Corollary~\ref{cor:homeo}, $\varphi:\Sigma_E^{\mathrm{irr}}\to(0,1]\setminus\mathbb Q$ is a homeomorphism; hence $\mathcal G_\phi$ is a dense $G_\delta$ in $(0,1]\setminus\mathbb Q$, and for $x\in\mathcal G_\phi$ with $\eta=f(x)$ we have
\[
\sum_{n=1}^\infty\phi\left(d_n(x)\right)=\sum_{n=1}^\infty\phi(\eta_n)=\infty.
\]
Since $(0,1]\setminus\mathbb Q$ is dense $G_\delta$ in $(0,1]$, it follows that $\mathcal G_\phi$ is a dense $G_\delta$ set in $(0,1]$. As $\mathcal G_\phi \subseteq \mathcal D_\phi$, we conclude that $\mathcal D_\phi$ is comeager (residual) in $(0,1]$.
\end{proof}

\begin{corollary}\label{cor:Engel-ls}
For $s>0$ put
\[
\mathcal D (s)\coloneqq\Bigl\{x\in(0,1]:\ \sum_{n=1}^{\infty}d_n(x)^{-s}=\infty\Bigr\}.
\]
Then $\mathcal D (s)$ contains a dense $G_\delta$ and is comeager in $(0,1]$.
\end{corollary}

\begin{proof}
Apply Theorem~\ref{thm:Gdelta} with $\phi(t)\coloneqq t^{-s}$.
\end{proof}

Recall the \emph{convergence exponent} of Engel digit sequences $\lambda:(0,1]\to[0,\infty]$ defined by
\[
\lambda(x)=\inf\left\{s\ge0: \sum_{n=1}^\infty d_n(x)^{-s}<\infty\right\}.
\]
For irrational $x$ it is classical that $\lambda(x)=\limsup_{n\to\infty}\frac{\log n}{\log d_n(x)}$ (see \cite{SW21}).

\begin{corollary}[{\cite[Theorem~3.5]{SW21}}]\label{cor:lambda-comeager}
The set $\{x\in(0,1]:\ \lambda(x)=\infty\}$ is comeager in $(0,1]$.
Equivalently, for every $\alpha\in[0,\infty)$, the set $\{x\in(0,1]:\ \lambda(x)\le \alpha\}$ is meager in $(0,1]$.
\end{corollary}

\begin{proof}
For each $k\in\mathbb N$, Corollary~\ref{cor:Engel-ls} gives that $\mathcal D (k)$ is comeager. Hence
\[
\{x\in(0,1]:\lambda(x)=\infty\}=\bigcap_{k=1}^\infty\mathcal D (k)
\]
is a countable intersection of comeager sets, thus comeager. For the equivalence, note that $\{x\in(0,1]:\sum_{n=1}^\infty d_n(x)^{-q}<\infty\}$ is meager for each rational $q>0$; then
\[
\{x\in(0,1]:\lambda(x)\le\alpha\}=\bigcup_{q\in\mathbb Q,\ q>\alpha}\left\{x\in(0,1]:\sum_{n=1}^\infty d_n(x)^{-q}<\infty\right\}
\]
is a countable union of meager sets, hence meager.
\end{proof}

\subsection{Concluding remarks}
The symbolic topology for Engel expansions yields a clean qualitative picture: $\varphi$ is globally Lipschitz, while $f$ is continuous exactly at irrationals and one–sided continuous at rationals. The comeager divergence phenomena in Theorem~\ref{thm:Gdelta} and Corollary~\ref{cor:Engel-ls} follow from the clopen cylinder structure on $\Sigma_E$ together with the homeomorphism $\varphi \colon \Sigma_E^{\mathrm{irr}}\to (0,1]\setminus\mathbb{Q}$; the fundamental intervals in $(0,1]$ are simply the mirror of this structure under $\varphi$. In short, the canonical Engel topology makes continuity and Baire category statements essentially tautological.

\end{document}